OpenStudy (anonymous):

Graph each pair of parametric equations. x=2t-1 y=t^2+5, -4

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

solve the first one for \(t\) then plug that in to the second one

OpenStudy (misty1212):

\[t=\frac{x+1}{2}\] plug that in for \(t\) in the second part, you will get a parabola

OpenStudy (anonymous):

wait so i dont even know how to solve for t..

OpenStudy (misty1212):

can you solve \[x=2t-1\] for \(t\) ?

OpenStudy (anonymous):

oh is it.. t=-1+x/2

OpenStudy (misty1212):

no, but close

OpenStudy (anonymous):

what did i do wrong?

OpenStudy (misty1212):

what did you do? takes two steps

OpenStudy (anonymous):

-2t=1+x t=(-1+x/2)

OpenStudy (misty1212):

\[x=2t-1\\ x+1=2t\\ \frac{x+1}{2}=t\]

OpenStudy (anonymous):

so i plug that into like desmos and that is my answer?

OpenStudy (loser66):

Why don't you make a T table?

OpenStudy (misty1212):

yeah like desmos what the monkey is "desmos"? and while i am asking, what the heck is a "T table"?

OpenStudy (anonymous):

i know.. what is a T tables haha

OpenStudy (anonymous):

desmos is a graphing calculator online

OpenStudy (loser66):